HETAV PATEL

Grade 11,

A vessel of volume 5 litres contains 1.4 g of N 2 and 0.4 g of He at 1500 K. If 30% of nitrogen molecules are dissociated into atoms then find the gas pressure.

A vessel of volume 5 litres contains 1.4 g of N2 and 0.4 g of He at 1500 K. If 30% of nitrogen molecules are dissociated into atoms then find the gas pressure.

Grade:12

1 Answers

Shobhit Varshney IIT Roorkee
9 years ago
Given total volume=5L
moles of nitrogen=1.4/14=0.1mol
30% dissociates into atoms
so 0.03 moles of N2 will convert into 0.06 mol of N
so totall number of moles of gaseous species=0.06(from N)+0.07mol(remainingN2)+ 0.1(He)=0.23
using ideal gas equation, PV=nRT:
.....................................P*5L=0.23*0.082*1500
...........now we can get P